OpenStudy (anonymous):

A field test for a new exam was given to randomly selected seniors. The exams were graded, and the sample mean and sample standard deviation were calculated. Based on the results, the exam creator claims that on the same exam, nine times out of ten, seniors will have an average score within 4% of 70%. Is the confidence interval at 90%, 95%, or 99%? What is the margin of error? Calculate the confidence interval and explain what it means in terms of the situation.

OpenStudy (anonymous):

@rational @robtobey @freckles

OpenStudy (anonymous):

@fallengoddess

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Stop tagging people and start showing your work. You can't have NOTHING.

OpenStudy (anonymous):

margin of error is 4%, the confidence interval would be 90% since says 9/10

OpenStudy (anonymous):

i dont know how to calculate confidence interval

OpenStudy (anonymous):

margin of error is 4% or 4% to 70%?

OpenStudy (tkhunny):

Margin of Error = \((z-score)\cdot (standard-deviation)\) Look up the z-scores for each of those confidence intervals. (2-tail)

OpenStudy (anonymous):

90 = 1.645 95 = 1.96 99 = 2.575

OpenStudy (anonymous):

So 1.645 - How do you find standard and deviation

OpenStudy (tkhunny):

That is awesome. Good work. You have to memorize it. Sample Proportion: Score: 0.70, Sample Standard Deviation: \(\sqrt{\dfrac{0.70\cdot(1 - 0.70)}{n}}\)

OpenStudy (anonymous):

I'm confused on what I'm supposed to do

OpenStudy (anonymous):

That problem is \[\sqrt{\frac{ 0.21 }{ n }}\]

OpenStudy (anonymous):

@perl

OpenStudy (perl):

ok im catching up here. you agree that the confidence level is 90%, since it says 9 out of 10

OpenStudy (anonymous):

yes

OpenStudy (perl):

this is a sample mean problem, not a sample proportion

OpenStudy (perl):

the margin of error is just 4 I believe

OpenStudy (anonymous):

i think so too

OpenStudy (perl):

so the confidence interval is ( 70 - 4, 70 + 4)

OpenStudy (perl):

We don't have enough information anyway to use that formula . the margin of error is already given as 4, since it says the sample mean is within 4 of 70. that means the sample mean ranges between 70 - 4, 70 + 4

OpenStudy (perl):

that is your confidence interval , at 90% confidence

OpenStudy (anonymous):

Well i understand what you wrote but what do i write for my answer?

OpenStudy (anonymous):

confidence interval = 90% margin of error = 4% but what do i write for calculate the confidence interval

OpenStudy (anonymous):

oh so just the 70 - 4 to 70 + 4 or 66 to 74

OpenStudy (perl):

confidence level 90% margin of error 4% The confidence interval is (66%, 74%) Interpretation: We are 90% confident that the true mean of the population is between 66% and 74% Another interpretation: If we were to take 100 additional samples, in 90 times the mean test score would fall between 66% and 74%.

OpenStudy (anonymous):

Ok I understand, thanks for all the help!!

OpenStudy (perl):

Your welcome :)

OpenStudy (tkhunny):

Whoops. Reading in my sleep, I guess.